# WUCG-25-05 YITP-25-63

### v1

@@ 2.1 Time diffeomorphism breaking
$\delta R_{\mu\nu\rho\sigma}\equiv R_{\mu\nu\rho\sigma}-2\frac{{\cal{H}}^{2}}{a^{2}}h_{\mu[\rho}h_{\sigma]\nu}+\frac{{\cal{H}}^{\prime}}{a^{2}}\left(h_{\mu\rho}n_{\nu}n_{\sigma}-h_{\mu\sigma}n_{\nu}n_{\rho}-(\mu\leftrightarrow\nu)\right)\,,$ (2.7)
@@ 2.2 Spatial symmetry breaking: $E$-$B$ basis
with the gauge coupling constant ${\cal{G}}$. Here, the rank-3 and rank-4 completely anti-symmetric symbols are denoted by $[abc]$ ($[123]=+1$) and $[\mu\nu\rho\sigma]$ ($[0123]=+1$), and the Levi-Civita tensor is $\epsilon_{\mu\nu\rho\sigma}=-\sqrt{-g}[\mu\nu\rho\sigma]$. Substituting Eq. (2.8) into the above definition, we obtain background electric and magnetic fields as
@@
At this point, it is worth clarifying why the symmetry-breaking $SU(2)_{I}\times SO(3)_{S}\to SO(3)_{D}$ will lead to parity violation. The key point is that the background (2.8) contains a non-zero magnetic component, which is odd under parity. As a result, the background is not invariant under the parity transformation. This should be contrasted with the case of a triplet of $U(1)$ gauge fields with the global $SU(2)$ symmetry [71, 72], which can be regarded as the $\mathcal{G}\to 0$ limit of our case. In the $U(1)$ case with the background (2.8), it has no magnetic field and parity is not violated through gauge field dynamics alone, since $U(1)$ lacks self-interactions. Parity violation can be induced, for instance, by a pseudoscalar field through the Chern-Simons coupling $\phi F\widetilde{F}$ [74, 16]. In the $SU(2)$ gauge fields case, on the other hand, we will see that the parity of GWs is inevitably broken regardless of whether there is a pseudoscalar field under the conditions we clarify later. We also note that our scenario is different from the solid inflation [60], where the residual symmetry of the solid is a global $SO(3)$ symmetry (and translations) rather than the local $SO(3)$ symmetry.
@@ 2.3 Alternative building blocks: $X$-$Y$ basis
$\delta E_{\mu\nu}$ $=\delta E^{(1)}_{\mu\nu}+\delta E^{(2)}_{\mu\nu}+\cdots,$ (2.28)
$\delta B_{\mu\nu}$ $=\delta B^{(1)}_{\mu\nu}+\delta B^{(2)}_{\mu\nu}+\cdots\,,$ (2.29)
@@
$\delta Y_{\mu\nu}$ $=\bar{B}\delta E^{(1)}_{\mu\nu}+\bar{E}\delta B^{(1)}_{\nu\mu}.$ (2.31)
@@
00 0 $=\frac{1}{2}(\delta E^{(1)}_{\mu\rho}\delta E^{(1)}_{\nu\sigma}h^{\rho\sigma}-\delta B^{(1)}_{\mu\rho}\delta B^{(1)}_{\nu\sigma}h^{\rho\sigma})+(\bar{E}\delta E^{(2)}_{(\mu\nu)}-\bar{B}\delta B^{(2)}_{(\mu\nu)}),$ (2.34)
00 0 $=\delta E^{(1)}_{\mu\rho}\delta B^{(1)}_{\nu\sigma}h^{\rho\sigma}+\bar{B}\delta E^{(2)}_{\mu\nu}+\bar{E}\delta B^{(2)}_{\nu\mu},$ (2.35)
@@
$\delta E^{(2)}_{\mu\nu}=\frac{-1}{2(\bar{E}^{2}+\bar{B}^{2})}\left[\bar{E}(\delta E^{(1)}_{\mu\rho}\delta E^{(1)}_{\nu\sigma}-\delta B^{(1)}_{\mu\rho}\delta B^{(1)}_{\nu\sigma})+2\bar{B}\delta E^{(1)}_{\rho(\mu}\delta B^{(1)}_{\nu)\sigma})\right]h^{\rho\sigma}\,,$ (2.36)
@@
[…]
@@ 3.1 Leading operators
As is well known, the dimensional counting in a non-trivial background can be different from that in a trivial background. The prominent example is the k-essence theory, where the Lagrangian is given by
$\mathcal{L}=\sum_{n=1}^{\infty}\frac{a_{n}}{M^{4(n-1)}}(\nabla\phi)^{2n}\,.$ (3.1)
The $n\geq 2$ operators are higher dimensional around the trivial background $\phi=0$ while they all can have the same importance as the $n=1$ operator if the background has a large velocity $\phi\sim M^{2}t$. Note that the EFT is still valid even around such a large background because higher-derivative operators like $\frac{1}{M^{2}}(\nabla^{2}\phi)^{2},\frac{1}{M^{6}}(\nabla\phi)^{2}(\nabla^{2}\phi)^{2}$, and $\frac{1}{M^{8}}(\nabla^{2}\phi)^{4}$ are well suppressed as long as the fluctuations $\delta\phi$ have a low-momentum $k\ll M$. Another example, especially relevant for us, is the chromo-natural inflation [29], which relies on the Chern-Simons coupling
@@
where $J^{\mu}$ is defined by $F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu}=\nabla_{\mu}J^{\mu}$. This higher dimensional operator becomes important if $\phi$ has a background velocity. The last example is the gauge-flation [28, 67] whose gauge part of the Lagrangian is given by
$-\frac{1}{4}F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu}+\frac{1}{16M^{4}}\left(F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu}\right)^{2}\,.$ (3.3)
Similarly to the k-essence, the operator $(F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu})^{2}$ becomes leading if the gauge field has a large background $\bar{E},\bar{B}\sim M^{2}$.
Our EFT aims to identify all the operators that can be leading in the broken phase and extract their universal features. In the language of $\phi$ and $F^{a}_{\mu\nu}$, the Lagrangian takes the form
@@
where $\cdots$ are operators containing more than one derivative of fields, e.g. $\nabla_{\mu}\nabla_{\nu}\phi$ and $\nabla_{\mu}F^{a}_{\nu\rho}$, which we ignore in the following. On the other hand, we keep all non-linear terms of first derivatives of the fields because, as we have seen, they can have the same importance in the broken phase. Note that we define our EFT in the Einstein frame in which the leading gravitational part of the Lagrangian takes the form of the Einstein-Hilbert action. [^5] If one wants to apply the EFT to dark energy, it may be useful to define the EFT in the Jordan frame where the matter fields are minimally coupled to gravity [75]. Using the building blocks developed in Sec. 2, we can write the same Lagrangian in a different basis:
@@
$\mathcal{L}=\frac{M_{\rm Pl}^{2}}{2}R+{\cal{L}}_{M}(h_{\mu\nu},\tau,\delta g^{00},\delta X_{\mu\nu},\delta Y_{\mu\nu})+\cdots\,.$ (3.6)
where $\delta K_{\mu\nu}$ and $\delta R_{\mu\nu\rho\sigma}$, which appear in the action of the EFT of inflation (2.4), are not included here because these are higher derivative operators [^6] Remembering the definitions of the unit vector (2.1) and the extrinsic curvature (2.3), we can see that $K_{\mu\nu}$ is related to a higher derivative operator $\nabla_{\mu}\nabla_{\nu}\phi$.. In this subsection, we adopt the $X$-$Y$ basis (3.6), and we will discuss how the Lagrangian is mapped into a different basis in Sec. 3.2.
@@ 3.2 Mapping between different bases
Now that we have explained the general procedure for mapping between bases, let us illustrate its application through five concrete examples. As we will see, all the operators appearing in our EFT action (3.8) can be generated by non-linear interactions in the $\phi$$F$ basis.
@@
    =¯+¯$=\bar{\cal{F}}+\bar{\cal{F}}_{X}\delta X+\bar{\cal{F}}_{Y}\delta Y+\frac{1}{2}\bar{\cal{F}}_{XX}\delta X^{2}+\bar{\cal{F}}_{XY}\delta X\delta Y+\frac{1}{2}\bar{\cal{F}}_{YY}\delta Y^{2}+\cdots$
    =¯¯$=\bar{\cal{F}}-\bar{\cal{F}}_{X}\bar{X}-\bar{\cal{F}}_{Y}\bar{Y}-\frac{1}{4}\bar{\cal{F}}_{X}F^{a}_{\mu\nu}F^{a\mu\nu}-\frac{1}{4}\bar{\cal{F}}_{Y}F_{\mu\nu}^{a}\widetilde{F}^{a\mu\nu}$
@@
      $\frac{1}{16}F^{a}_{\mu\nu}F^{b\mu\nu}F^{a}_{\rho\sigma}F^{b\rho\sigma}=-\frac{1}{3}\bar{X}^{2}-\frac{1}{6}\bar{X}F^{a}_{\mu\nu}F^{a\mu\nu}+\delta X_{\mu\nu}\delta X^{\mu\nu}+\cdots\,.$ (3.22)
    • $[abc]F^{a}F^{b}F^{c}$
@@
      Here we used the relation ${[abc]e^{a}_{\mu}e^{b}_{\nu}e^{c}_{\rho}=\epsilon_{\mu\nu\rho}}$.
@@
    $=\bar{\phi}^{\prime}{}^{2}\left[-\frac{3\bar{E}^{2}}{2a^{2}}+\frac{3}{2}\bar{E}^{2}\delta g^{00}-\frac{\bar{E}}{a^{2}}\delta E+\bar{E}\delta g^{00}\delta E-\frac{1}{2a^{2}}\delta E_{\mu\nu}\delta E^{\mu\nu}+\cdots\right]$

### v2

@@ 2.1 Time diffeomorphism breaking
$\delta R_{\mu\nu\rho\sigma}\equiv R_{\mu\nu\rho\sigma}-2\frac{{\cal{H}}^{2}}{a^{2}}h_{\mu[\rho}h_{\sigma]\nu}+\frac{{\cal{H}}^{\prime}}{a^{2}}\left(h_{\mu\rho}n_{\nu}n_{\sigma}-h_{\mu\sigma}n_{\nu}n_{\rho}-(\mu\leftrightarrow\nu)\right)\,.$ (2.7)
@@ 2.2 Spatial symmetry breaking: $E$-$B$ basis
with the gauge coupling constant $\cal{G}$}. Here, the rank-3 and rank-4 completely anti-symmetric symbols are denoted by $[abc]$ ($[123]=+1$) and $\mu\nu\rho\sigma]$ ($[0123]=+1$), and the Levi-Civita tensor is $\epsilon_{\mu\nu\rho\sigma}=-\sqrt{-g}[\mu\nu\rho\sigma]$. Substituting Eq. (2.8) into the above definition, we obtain background electric and magnetic fields as
@@
At this point, it is worth clarifying why the symmetry-breaking $SU(2)_{I}\times SO(3)_{S}\to SO(3)_{D}$ will lead to parity violation. The key point is that the background (2.8) contains a non-zero magnetic component, which is odd under parity. As a result, the background is not invariant under the parity transformation. This should be contrasted with the case of a triplet of $U(1)$ gauge fields with the global $SU(2)$ symmetry [71, 72], which can be regarded as the $\mathcal{G}\to 0$ limit of our case. In the $U(1)$ case with the background (2.8), it has no magnetic field and parity is not violated through gauge field dynamics alone, since $U(1)$ lacks self-interactions. Parity violation can be induced, for instance, by a pseudoscalar field through the Chern-Simons coupling $\phi F\widetilde{F}$ [74, 16]. In the $SU(2)$ gauge fields case, on the other hand, we will see that the parity of GWs is inevitably broken regardless of whether there is a pseudoscalar field under the conditions we clarify later.
Let us make a comparison with the solid inflation where the spatial symmetry is also broken [60]. The solid inflation is based on a triplet of scalar fields, $\phi^{I}$ ($I=1,2,3$), with the global internal rotation and translation symmetries, $\phi^{I}\rightarrow O^{I}{}_{J}\phi^{J}$ ($O^{I}{}_{J}\subset SO(3)$), and $\phi^{I}\rightarrow\phi^{I}+a^{I}$ ($a^{I}=\text{const.}$). The scalar fields are assumed to have the background configuration $\langle\phi^{I}\rangle=x^{I}$, which leads to the spontaneous symmetry breaking $SO(3)_{I}\times SO(3)_{S}\to SO(3)_{D}$ (the translation is also broken), similarly to (2.16). However, the residual symmetries of the models with gauge fields and the solid inflation are different: the former possesses the local $SO(3)_{D}$ symmetry while the latter only has the global $SO(3)_{D}$ symmetry. In particular, these models have different numbers of degrees of freedom in the tensor modes. The gauge field models have four tensor modes in the broken phase, where two originate from the metric and other two are from the gauge fields. On the other hand, the solid inflation only has two tensor modes from the metric, acquiring a mass due to the symmetry breaking.
@@ 2.3 Alternative building blocks: $X$-$Y$ basis
$\delta E_{\mu\nu}$ $=\delta E^{(1)}_{\mu\nu}+\delta E^{(2)}_{\mu\nu}$$\cdots$, (2.28)
$\delta B_{\mu\nu}$ $=\delta B^{(1)}_{\mu\nu}+\delta B^{(2)}_{\mu\nu}$$\cdots\,$, (2.29)
@@
$\delta Y_{\mu\nu}$ $=\bar{B}\delta E^{(1)}_{\mu\nu}+\bar{E}\delta B^{(1)}_{\nu\mu}$, (2.31)
@@
$0$ $=\frac{1}{2}(\delta E^{(1)}_{\mu\rho}\delta E^{(1)}_{\nu\sigma}h^{\rho\sigma}-\delta B^{(1)}_{\mu\rho}\delta B^{(1)}_{\nu\sigma}h^{\rho\sigma})+(\bar{E}\delta E^{(2)}_{(\mu\nu)}-\bar{B}\delta B^{(2)}_{(\mu\nu)}),$ (2.34)
$0$ $=\delta E^{(1)}_{\mu\rho}\delta B^{(1)}_{\nu\sigma}h^{\rho\sigma}+\bar{B}\delta E^{(2)}_{\mu\nu}+\bar{E}\delta B^{(2)}_{\nu\mu}$, (2.35)
@@
$\delta E^{(2)}_{\mu\nu}=\frac{-1}{2(\bar{E}^{2}+\bar{B}^{2})}\left[\bar{E}(\delta E^{(1)}_{\mu\rho}\delta E^{(1)}_{\nu\sigma}-\delta B^{(1)}_{\mu\rho}\delta B^{(1)}_{\nu\sigma})+2\bar{B}\delta E^{(1)}_{\rho(\mu}\delta B^{(1)}_{\nu)\sigma}\right]h^{\rho\sigma}\,$, (2.36)
@@
By substituting (2.32) and (2.33) into them, we obtain the inverse formula up to the second order in perturbations. It is cumbersome but straightforward to find higher-order solutions for $\delta E_{\mu\nu}$ and $\delta B_{\mu\nu}$ by repeating these procedures provided $\bar{E}\neq 0$. [^4] The $X$-$Y$ basis itself can be used even for $\bar{E}=0$ if $\bar{B}\neq 0$. In this case, one can use the gauge condition $\delta_{aA}B^{a}_{[\mu}e^{A}_{\nu]}=0$ to define the new basis and show the equivalence of this to the $X$-$Y$ basis. If necessary, $\delta E^{a}_{\mu}$ and $\delta B^{a}_{\mu}$ are obtained by applying (2.21).
@@ 3.1 Leading operators
As is well known, the dimensional counting in a non-trivial background can be different from that in a trivial background. Let us see some explicit examples in the flat spacetime. The prominent example is a k-essence scalar field, where the Lagrangian is given by
$\mathcal{L}=\sum_{n=1}^{\infty}\frac{a_{n}}{M^{4(n-1)}}(\nabla\phi)^{2n}\,$, (3.1)
with $M$ being a constant of mass dimension one. The $n\geq 2$ operators are sub-leading around the trivial background $\bar{\phi}=0$. On the other hand, they all can have the same importance as the $n=1$ operator if the background has a large velocity $\bar{\phi}\sim M^{2}t$. Substituting $\phi=M^{2}t+\delta\phi$ into the Lagrangian, we obtain
$(n=1),\quad(\nabla\phi)^{2}\supset(\nabla\delta\phi)^{2}\sim k^{2}\delta\phi^{2}\,$, (3.2)
$(n\geq 2),\quad\frac{1}{M^{4(n-1)}}(\nabla\phi)^{2n}\supset\frac{1}{M^{4(n-1)}}(\nabla\bar{\phi})^{2(n-1)}(\nabla\delta\phi)^{2}\sim k^{2}\delta\phi^{2}\,.$ (3.3)
Here, only leading perturbations are shown in each operator and we assume that $\delta\phi$ fluctuates with a momentum $k$. We can see that $n\geq 2$ operators lead to the same order as $n=1$ operator due to the non-trivial background with the large velocity $\bar{\phi}=M^{2}t$. Note that the EFT is still valid even around such a large background because higher-derivative operators are well suppressed as long as the fluctuations $\delta\phi$ have a low-momentum $k\ll M$. For instance,
$\frac{1}{M^{2}}(\nabla^{2}\phi)^{2}\supset\frac{1}{M^{2}}(\nabla^{2}\delta\phi)^{2}\sim\frac{k^{4}}{M^{2}}\delta\phi^{2}\,$, (3.4)
$\frac{1}{M^{6}}(\nabla\phi)^{2}(\nabla^{2}\phi)^{2}\supset\frac{1}{M^{6}}(\nabla\bar{\phi})^{2}(\nabla^{2}\delta\phi)^{2}\sim\frac{k^{4}}{M^{2}}\delta\phi^{2}\,$, (3.5)
which are suppressed by $k^{2}/M^{2}\ll 1$ compared to the leading operators (3.1). In the following, we only focus on quadratic terms in perturbations and count the number of the highest-order derivatives acting on perturbations. Then, the discussions so far can be summarised as
$\frac{1}{M^{4(n-1)}}(\nabla\phi)^{2n}\sim k^{2}\,,\qquad\frac{1}{M^{4n+2}}(\nabla\phi)^{2n}(\nabla^{2}\phi)^{2n}\sim\frac{k^{4}}{M^{2}}\ll k^{2}\,.$ (3.6)
The first one is leading, whereas the second one is sub-leading within the regime of validity of the EFT, $k\ll M$.
Let us now apply the same counting rule of derivatives to concrete examples. Another example, especially relevant for us, is the chromo-natural inflation [29], which relies on the Chern-Simons coupling
@@
where $J^{\mu}$ is defined by $F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu}=\nabla_{\mu}J^{\mu}$, which only yields up to the first derivative of perturbations. Therefore, if $\phi$ has a background velocity, we can count the number of derivatives as
$\frac{\lambda}{f}\phi\,F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu}\sim\frac{\lambda M^{2}}{f}k\,.$ (3.8)
This operator is thus more important than the kinetic term $FF\sim k^{2}$ in the low-energy regime $\lambda M^{2}/f\gg k$. The other example is the gauge-flation [28, 67] whose gauge part of the Lagrangian is given by
${\cal{L}}=-\frac{1}{4}F^{a}_{\mu\nu}{F}^{a\mu\nu}+\frac{1}{16M^{4}}\left(F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu}\right)^{2}\,.$ (3.9)
Similarly to the k-essence, the operator $(F^{a}_{\mu\nu}\widetilde{F}^{a\mu\nu})^{2}$ becomes leading if the gauge field has a large background $\bar{E},\bar{B}\sim M^{2}$, that is, $\frac{1}{M^{4}}F^{4}\sim k^{2}$. On the other hand, operators involving derivatives of the field strength are sub-leading as they yield higher derivatives, e.g., $\frac{1}{M^{2}}(\nabla F)^{2}\sim\frac{k^{4}}{M^{2}}$.
Gravitational couplings can be counted similarly. The couplings $f(\phi)R$ involve only up to second derivatives, and thus can be leading operators. The gravitational Chern-Simons coupling we mentioned at the beginning of Sec. 2 is
$\frac{\phi}{f}\epsilon^{\mu\nu\rho\sigma}R_{\alpha\beta\mu\nu}R^{\alpha\beta}{}_{\rho\sigma}=-\frac{1}{f}\nabla_{\mu}\phi\,\mathcal{J}^{\mu}+(\text{boundary terms})\sim\frac{M^{2}}{fM_{\rm Pl}^{2}}k^{3}$ (3.10)
where we have first performed the integration by parts and used $\bar{\phi}=M^{2}t$. The Planck mass $M_{\rm Pl}$ appears after canonical normalisation of metric perturbations. Hence, the gravitational Chern-Simons coupling is third order in derivatives, and thus sub-leading. One can also discuss derivative couplings to the curvature, such as $R(\nabla\phi)^{2}$ and $RF^{2}$, which are higher derivatives and should be suppressed, in general.
All in all, the leading operators here refer to leading in the sense of derivative expansion, rather than counting mass dimensions. These operators describe low-energy/momentum dynamics of perturbations around a non-trivial background. Note that there are exceptions to the discussion thus far. It is well known, in the context of modified gravity like Horndeski theories [75], that certain combinations of higher derivative operators generate equations of motion up to second order in derivatives. These combinations can be counted as leading by following the rule of the number of derivatives, albeit requiring two mass scales often denoted by $\Lambda_{2}$ and $\Lambda_{3}$ with the hierarchy $\Lambda_{2}\gg\Lambda_{3}$ [76]. For the sake of simplicity, this paper will consider the conservative case where no such hierarchy exists, [^5] It has been shown in [77, 78, 79] that this hierarchy obstructs UV completion (or renders a low cutoff) at least around the Minkowski background. even though the framework of our EFT can easily extend to accommodate such combinations.
Under these rules, let us identify all leading operators and extract their universal features. In the language of $\phi$ and $F^{a}_{\mu\nu}$, the most general Lagrangian takes the form
@@
where $\cdots$ are operators containing more than one derivative of fields, e.g. $\nabla_{\mu}\nabla_{\nu}\phi,\nabla_{\mu}F^{a}_{\nu\rho}$ and derivative couplings to the curvature, which we ignore in the following. On the other hand, we keep all non-linear terms of first derivatives of the fields because, as we have discussed, they can have the same importance in the broken phase. We define our EFT in the Einstein frame by eliminating the coupling $f(\phi)R$ through a field redefinition so that the leading gravitational part of the Lagrangian takes the form of the Einstein-Hilbert action. [^6] When applying the EFT to dark energy, it may be useful to define the EFT in the Jordan frame where the ordinary matter fields are kept minimally coupled to gravity [80]. Using the building blocks developed in Sec. 2, we can write the same Lagrangian in a different basis:
@@
$\mathcal{L}=\frac{M_{\rm Pl}^{2}}{2}R+{\cal{L}}_{M}(h_{\mu\nu},\tau,\delta g^{00},\delta X_{\mu\nu},\delta Y_{\mu\nu})+\cdots\,$, (3.13)
[…]
@@ 3.2 Mapping between different bases
[…]
@@
    $=\bar{\cal{F}}+\bar{\cal{F}}_{X}\delta X+\bar{\cal{F}}_{Y}\delta Y+\frac{1}{2}\bar{\cal{F}}_{XX}\delta X^{2}+\bar{\cal{F}}_{XY}\delta X\delta Y+\frac{1}{2}\bar{\cal{F}}_{YY}\delta Y^{2}+\cdots$
    $=\bar{\cal{F}}-\bar{\cal{F}}_{X}\bar{X}-\bar{\cal{F}}_{Y}\bar{Y}-\frac{1}{4}\bar{\cal{F}}_{X}F^{a}_{\mu\nu}F^{a\mu\nu}-\frac{1}{4}\bar{\cal{F}}_{Y}F_{\mu\nu}^{a}\widetilde{F}^{a\mu\nu}$
@@
      $\hskip-14.22636pt\frac{1}{16}F^{a}_{\mu\nu}F^{b\mu\nu}F^{a}_{\rho\sigma}F^{b\rho\sigma}=-\frac{1}{3}\bar{X}^{2}-\frac{1}{6}\bar{X}F^{a}_{\mu\nu}F^{a\mu\nu}+\delta X_{\mu\nu}\delta X^{\mu\nu}+\cdots\,.$ (3.29)
    • $abc]F^{a}F^{b}F^{c}$
@@
      Here we used the relation $abc]e^{a}_{\mu}e^{b}_{\nu}e^{c}_{\rho}=\epsilon_{\mu\nu\rho}$}.
@@
    $=\bar{\phi}^{\prime}{}^{2}\left[-\frac{3\bar{E}^{2}}{2a^{2}}+\frac{3}{2}\bar{E}^{2}\delta g^{00}-\frac{\bar{E}}{a^{2}}\delta E+\bar{E}\delta g^{00}\delta E-\frac{1}{2a^{2}}\delta E_{\mu\nu}\delta E^{\mu\nu}$$\cdots\right$]